OpenStudy (anonymous):

Write an equation of a line in slope intercept form with slope -3 going through the point (2, - 2).

OpenStudy (anonymous):

Use point-slope form: y - y1 = m(x - x1) your slope = m = -3 your point (x1, y1) = (2, -2) Just plug that in and solve for y to get your answer. Show your own freaking work.

OpenStudy (anonymous):

Thank You So Much!

OpenStudy (anonymous):

This was found in your Lessons and I know you can reach out to your teacher for help so please take advantage of them instead of asking help on here when you copy/paste a question. The better thing to say would have been this: "Hey, I don't understand how to write the equation of a line with a point and the slope, any tips?"